OpenStudy (anonymous):

SOMEBODY QUIZ ME ON Factoring Trinomials of the Form x2 + bx + c and x2 - bx + c please quiz tomorrow need to see if i understand

OpenStudy (carson889):

Factorize these: x^2 -3x - 4=0 x^2 + 4x + 4 =0 2x^2 -7x -15 =0 Bonus: x^3 + 3x^2 -18x = 0

OpenStudy (anonymous):

1) (x + 1) (x -4) 2) (x + 2) (x + 2) 3) i need help one and the bonus ! @carson889

OpenStudy (anonymous):

help on 3 i meant !

OpenStudy (carson889):

We will use a trick to do this: